if a and b are two vectors such that |a vector+b vector|=root 29 and a vector cros(2i^+4j^+4k^)=(2i^+4j^+4k^)cross b vector then find the value of (a vector+b vector).(-7i^+2j^+3k^)

Grade:

1 Answers

Arun
5 years ago
Dear student
 
a x (2, 4, 4) + b x (2,4, 4) = 0
 
Or we can write
 
(a+b )  x (2 i + 4j + 4k) = 0
 
Hence
 
a + b = p( 2i + 4j + 4k)
 
Now
 
As |a +b| = sqrt (29)
 
4p² + 16p² + 16p² = 29
 
p² = 29/36
 
Hence
 
p = sqrt(29) /6
 
Now
 
You can calulate the further question.